### v2

# Toward Intelligent and Secure Cloud: Large Language Model Empowered Proactive Defense

Abstract
The rapid evolution of cloud computing technologies and the increasing number of cloud applications have provided numerous benefits in our daily lives. However, the diversity and complexity of different components pose a significant challenge to cloud security, especially when dealing with sophisticated and advanced cyberattacks. Recent advancements in generative foundation models (GFMs), particularly in the large language models (LLMs), offer promising solutions for security intelligence. By exploiting the powerful capabilities in language understanding, data analysis, task inference, action planning, and code generation, we present LLM-PD, a novel defense architecture that proactively defeats various threats. LLM-PD can efficiently make decisions through comprehensive data analysis and sequential reasoning, as well as dynamically create and deploy actionable defense mechanisms on the target cloud. Furthermore, it can flexibly self-evolve based on experience learned from previous interactions and adapt to new attack scenarios without additional training. The experimental results demonstrate its remarkable ability in terms of defense effectiveness and efficiency, particularly highlighting an outstanding success rate when compared with other existing methods.
Index Terms:
Cloud Computing, Large Language Models, Proactive Defense, Security Intelligence, Self Evolution.
I Introduction
Cloud computing has accelerated rapidly in recent years, establishing itself as a significant paradigm in service-oriented computing. In the last decade, cloud computing has received significant attention due to its notable advantages such as scalability, flexibility, and increased cost-efficiency. For example, cloud computing has been integrated into the daily lives of businesses and individuals, offering a diverse range of services (e.g., storage, networking, and platform provisions) seamlessly accessible from any geographical location connected to the Internet.
Despite the obvious benefits of cloud computing, the diverse and complex nature of its various components, i.e., network, architecture, APIs, and hardware, has led to heightened security apprehensions. Due to the utilization of standard Internet protocols and virtualization methods, it is susceptible to potential security breaches. These vulnerabilities may stem from conventional sources such as IP spoofing, Address Resolution Protocol spoofing, and Distributed Denial of Service (DDoS) attacks, among others. Additionally, emerging threats such as zero-day attacks, characterized by their unknown nature, pose a considerable challenge to traditional solutions, which may not be sufficiently effective in mitigating such attacks.
To address the aforementioned challenges, several proactive defense techniques have been proposed, including Moving Target Defense (MTD) [1], cyber deception [2], Mimic Defense [3], among others. These methods emphasize the proactive identification, warning, and response to potential threats through automated and adaptive mechanisms, either before or during an attack, thereby effectively reducing security risks and minimizing potential losses.
Although these solutions overcome, to some extent, the shortcomings of traditional solutions, they require modifications to mitigation mechanisms that may not be effective across diverse environments. Moreover, the decision-making of defense deployment predominantly relies on heuristic, machine learning, and deep learning algorithms. Nevertheless, given the increasing complexity of cloud-based applications and the wide array of attack vectors, it is hard for any specific strategy to fully adapt to the different and time-varying scenarios in the cloud. For this reason, there is an urgent requirement for an intelligent and adaptable approach to facilitate proactive defense within the cloud environment.
Fortunately, as a typical example of Generative Foundation Models (GFMs) that can be adapted for specific purposes, Large Language Models (LLMs) have made a profound impact on research and practical applications [4]. Specifically, LLMs excel in addressing general tasks through in-context learning, generating multimodal data via prompt engineering, breaking down complex problems into simpler components through strategic reasoning, simulating specified scenarios through role-playing, and making decisions based on prior knowledge and situational understanding [5]. Building upon these advantages, recent research has leveraged LLMs in the cybersecurity community, empowering security professionals to explore various attack vectors (e.g., vulnerability detection) and develop autonomous agents (e.g., code fixing). Therefore, the extensive knowledge encoded in LLMs has sparked our interest in exploring their potential for comprehensive protection in the complex and ever-changing cloud security scenarios.
In this paper, differing significantly from the previous defense using expert knowledge or specific strategies, we leverage pre-trained LLMs with different prompts and deploy them across a variety of attack scenarios in the cloud domain, achieving enhanced protection and proactive response. The main contribution of this work includes the following.
  • We design a novel robust and efficient cloud security architecture called LLM-PD. To the best of our knowledge, this is the first comprehensive end-to-end LLM-driven proactive defense architecture where specialized agents handle different aspects of the defense lifecycle while maintaining contextual awareness.
  • We achieve cross-attack adaptability without extensive training, overcoming limitations of traditional ML methods that require retraining for different scenarios. Moreover, we develop a self-evolving memory feedback mechanism that enables continuous learning and improvement of defense capabilities through experience.
  • We conduct a case study with GPT-4o mini to demonstrate practical implementation of our architecture and its adaptation to various DoS attacks. Comparative experiments reveal significant advantages in surviving rate, time efficiency, and success rate over existing approaches. The source code is available at https://github.com/SEU-ProactiveSecurity-Group/LLM-PD.
The rest of this paper is organized as follows. We commence with the background of proactive defense and discuss the related work of LLM for cybersecurity. Then we elaborate on the proposed architecture with design details introduced. Next, we discuss the prototype construction and use the mitigation of multiple DoS attacks as a case study. We conduct experimental evaluations to illustrate efficient performance. Finally, we conclude the work and analyze future challenges.
II Background and Related Work
II-A Proactive Defense in Cloud Networks
Proactive defense empowers cybersecurity personnel with anticipatory control over system defenses, establishing a tactical advantage and elevating protection to a more sophisticated level. Unlike the reactive approach, it involves real-time monitoring of the system, enabling the rapid detection of changes in network traffic, analysis of program behavior, and the prevention of suspicious activities to ensure security.
There have been some efforts that attempted to enhance security in a proactive manner. For example, Zhou et al. [6] seamlessly combine MTD technologies and Deep Reinforcement Learning (DRL) to proactively defend against Low-rate DDoS attacks. This approach introduces a Markov decision process (MDP) to formalize attack scenarios and adopts a Deep Q-network (DQN) algorithm to mitigate the impact of attacks in a resource-efficient manner.
Ma et al. [7] leverage the mutation-enabled defense mechanism in the cloud-native applications for mitigating Man-in-The-Middle (MITM) attacks. Driven by a maximizing coloring-based mutation grouping algorithm, it improves the robustness against attacks while guaranteeing the continuity of network connection with an acceptable performance loss.
Wu et al. [8] propose a security defense framework based on Intrinsic Cloud Security (ICS). It fuses the MTD concept and mimic defense theory within the Network Functions Virtualization (NFV) based cloud and subsequently implements the integration to achieve proactive protection against multiple threats, including cache side-channel attacks, code reuse attacks, and latent persistent threats.
However, existing work lacks intelligent decision-making and typically relies on predefined rules or limited policy sets that cannot adapt to complex and evolving threats. Although some methods incorporate intelligent algorithms, they require specialized training over thousands of episodes for each attack type and do not generalize well to even slightly different scenarios without extensive retraining. In contrast, LLM-PD offers cross-scenario defense capabilities without requiring retraining for different attack scenarios, and can dynamically generate and deploy entirely new defense measures.
II-B LLM for Cybersecurity Enhancement
Traditional approaches, such as signature-based detection and rule-based systems, often struggle to keep pace with the ever-changing threat landscape in the dynamic cyberspace. However, advancements in LLMs have marked a significant shift in cybersecurity paradigms, offering more adaptive and intelligent technologies [9]. Specifically, the functionalities of LLMs have contributed to promising enhancements in defense applications, including threat detection, vulnerability analysis, automated defense mechanisms, and others.
For instance, SecurityBERT [10] processes network traffic by integrating the Privacy-Preserving Fixed-Length Encoding (PPFLE) method with the Byte-level Byte-Pair Encoder (BBPE) Tokenizer. It employs the Bidirectional Encoder Representations from Transformers (BERT) model in identifying fourteen types of attacks with high accuracy.
Furthermore, Yin et al. [11] propose a pipeline for quantitatively evaluating LLMs’ performance in various software vulnerability-related tasks using publicly available datasets. Their evaluation results indicate that while LLMs perform well in certain aspects, there is still a need for improvement in understanding the subtle differences in code vulnerabilities and in their ability to describe these vulnerabilities, in order to fully realize their potential.
Additionally, an LLM-based software honeypot named SheLLM [12] was introduced to generate Linux-like shell output. This honeypot aims to evaluate the model’s credibility by having security experts assess its responses to attackers’ commands. The experimental results suggest that SheLLM can produce credible and dynamic responses, effectively addressing the limitations of existing honeypots.
Existing LLM applications in security primarily focus on isolated security functions and cannot ensure lifecycle protection over a relatively long time period. On the contrary, we integrate complete defense lifecycle into a cohesive end-to-end intelligent system and realize continuous self-optimization through memory-based iterative learning.
III LLM Empowered Proactive Defense Architecture for Cloud Security
To achieve an autonomous, AI-driven, and secure-by-design defense implementation, we present an architecture, named LLM-PD, that deeply integrates LLMs with existing security features, as illustrated in Fig. 1. LLM-PD goes beyond simple prompt engineering by introducing a sophisticated, security-specific multi-agent system designed for cloud environments. Each component incorporates a specific LLM-driven agent, all of which together build up a comprehensive proactive defense pipeline.
Figure 1: An overview of the proposed LLM empowered proactive defense architecture in the cloud networks.
III-A Data Collection and Reconstruction
In cloud networks, the extensive deployment of network applications results in a substantial volume of data available for security purposes. This data includes network traffic, performance metrics, logs, events, and service configurations, etc. Essentially, each type of data is considered independent and is formatted differently. For example, Wireshark is a tool that used to capture network traffic, typically in .pcap format, whereas alarm events are generated in .txt format based on firewall rules. Therefore, the ability to proficiently use multiple security tools and analyze massive heterogeneous data poses a significant challenge, even for cybersecurity professionals.
Powered by the extraordinary prowess of LLMs, this component works as a collector, which is generally fed with a prompt of natural language descriptions that cover the range of targets, tools, time, and reconstruction format. In detail, the collector is first designed to call multiple security tools to gather data from the target cloud network within given time period. Specifically, this collector can adapt to new collection tools through natural language interactions. Different from typical approaches, our collector implements semantically-driven heterogeneous data integration that can process unstructured security data from multiple sources. Nevertheless, given the large number of data sources, redundant and irrelevant information is inevitable. To this end, the collector aggregates data by evaluating the relationships between them and eliminating duplicate contents. Finally, this module reconstructs the refined results into a standardized representation (e.g., JSON file) and transmits it to the next stage for further processing.
III-B Status and Risk Assessment
Upon receiving the formatted information from the collector, we necessitate another module to serve as an analyzer to perform a thorough assessment of system status and identify potential security risks.
III-B1 Status Analysis
Within the status analysis function, information regarding hardware (e.g., power consumption), system (e.g., system load), network (e.g., inflow and outflow traffic volumes), and applications (e.g., number of connections) can be readily extracted from specific fields in the received standardized file. Note that the status information of the target cloud network not only reflects the operational state of the system but also allows for the derivation of constraints for subsequent defense tasks. For example, when the system utilization is high, it implies that the defense objectives must be achieved under conditions of limited available resources.
III-B2 Risk Evaluation
The risk evaluation function primarily focuses on the security indicators of cloud networks. This function utilizes information collected from various monitoring tools, in conjunction with historical normal data, to identify potential threats and anomalies within the system. Concurrently, the risk level is quantified (e.g., on a scale of 0-10) based on the scope, impact, and duration of the threat, thereby prioritizing subsequent defense tasks. Additionally, in cases of cloud network failures caused by severe attacks or malicious disruptions (such as service interruptions), this feature can also leverage existing information for fault localization, effectively identifying attack targets and generating defense goals.
Unlike rule-based analyzers, we implement multi-level threat modeling and quantitative risk assessment combined with causal reasoning to understand the relationships between different attack indicators and system anomalies.
III-C Task Inference and Decision-Making
Based on the evaluation results from the analyzer, this module, functioning as a decision-maker, infers the defense tasks that need to be completed and formulates corresponding strategies. In complex cloud environments, however, multiple threats or failure points may arise simultaneously, requiring the simultaneous achievement of multiple defense objectives within a single overarching task. To prevent task conflicts and facilitate efficient decision-making, we have developed a hierarchical agent based on LLMs that decomposes complicated defense tasks into multiple subtasks, enabling the independent development of defense policies for each subtask.
III-C1 Task Decomposition
In a high-level threat scenario involving multiple defense objectives, it is crucial to clearly plan and generate detailed tasks for better strategic reasoning. To this end, we have designed a task decomposition function that breaks down complex tasks into multiple independent subtasks. We require the LLM to delineate the implementation constraints of each subtask, assign priorities based on the previously mentioned risk levels, and analyze the dependencies among tasks for sequential arrangement.
III-C2 Inference and Decision-Making
In this process, tasks are executed sequentially according to their assigned order. For each subtask, this function systematically deduces the necessary step-by-step process required for decision-making, with the logic outlined as follows. First, details of the task are extracted from each subtask. Then, based on the current task requirements and constraints, human-like sequential reasoning for defensive actions over future attack-defense interactions is conducted. Additionally, this function is designed not only to enhance security but also to incorporate preferences, such as focusing on reducing defense overhead or ensuring user experience. These preferences can be preset by security personnel and dynamically adjusted during implementation. Finally, by balancing multiple competing objectives, it imports solvers to optimize the policy among the available strategies.
III-D Defense Deployment and Execution
This module, which plays the role of deployer, is responsible for the deployment, implementation, and execution of defense strategies. Specifically, upon receiving a defense strategy, the module first extracts the defense mechanisms and actions included in the strategy. Then, it queries these defense measures involved to the matching function and determines if there are any corresponding items with the existing defense database.
When a match is successful, the deployer proceeds to extract additional detailed information from the strategy, including deployment targets, execution commands, and implementation parameters. Subsequently, the module leverages these configurations to invoke the existing defense actions, ensuring that they are aligned with the strategic objectives. These actions are then executed automatically within the target cloud network, facilitating efficient deployment of defense mechanisms without the need for manual intervention.
When a particular mechanism does not exist in the defense library, the deployer harnesses the capabilities of LLMs to invoke programming languages, thereby generating useful scripts based on carefully crafted prompts to construct the necessary defense mechanism. Once these scripts are generated, they undergo a compilation and verification process to ensure their reliability and operational compatibility, addressing a critical limitation in automated defense generation that distinguishes our approach from existing work. Upon successful verification, the scripts are deployed and executed within the cloud infrastructure, effectively implementing the desired defense mechanism according to the generated policy. Simultaneously, these newly generated scripts are archived in the defense library, enriching the repository with new capabilities and ensuring that similar requirements in the future can be addressed more efficiently.
III-E Effectiveness Analysis and Feedback
After the execution of defense actions, this module plays the role of feedback giver to determine whether the current threat has been mitigated and to offer insights for future attack-defense interactions. By conducting a comprehensive evaluation of the defense effectiveness, the module quantifies the performance of the defense mechanisms post-execution, verifies if it meets expectations, and subsequently generates feedback for the LLM to incorporate into its memory.
III-E1 Multi-dimensional Evaluation
The evaluation of defense effectiveness in LLM-PD employs a multi-dimensional assessment approach that goes beyond analyzing a single security metric. Our framework integrates security effectiveness, recovery time, resource consumption, financial cost, and Quality of Service (QoS) to achieve a more comprehensive and rational assessment of effectiveness. Each dimension is assigned an importance weight based on the specific defense context and organizational priorities. This quantitative effectiveness score enables the LLM agent to make nuanced decisions that balance multiple competing objectives.
III-E2 Endpoint Validation
The primary purpose of the endpoint validation function is to determine whether the current defense process has reached its endpoint. After each defensive action is deployed, this function evaluates the current system state against success mitigation and failure conditions. When the validation function determines that the defense round has concluded, it tags the complete sequence of defensive actions implemented during the current episode with outcome flags (success or failure) and relevant contextual metadata. The annotated defense sequence is then transmitted to the memory module for storage and future reference.
III-E3 Memory-Based Optimization
The memory module implements a sophisticated dual-layer memory architecture that consists of intra-episode pool and inter-episode pool. Within each defense episode, the intra-episode memory pool maintains action-effectiveness sequences that associate each defensive action with quantitative effectiveness metrics. These action-effectiveness pairs guide the LLM agent toward selecting progressively higher-effectiveness defensive actions within the ongoing episode. For the inter-episode memory pool, it implements a sliding window mechanism that stores the most recent batches of validation results, that is, complete defense action sequences along with their overall outcomes. This module addresses token limit constraints while preserving the most relevant historical knowledge, prioritizing successful action sequences that required the fewest steps and avoiding previously failed defense actions. This iterative feedback loop enables a dynamic learning process and allows the system to adapt and refine its strategies over time, enhancing its resilience and effectiveness in the face of evolving security challenges. Importantly, this entire process occurs without the need for human intervention or expert knowledge, facilitating continuous improvement across multiple defense episodes.
III-F Practical Applications and Deployment Scenarios
The LLM-PD architecture can be effectively deployed across various cloud security scenarios. We highlight three critical application areas as follows.
III-F1 Mitigating DoS/DDoS Attacks in Cloud Services
LLM-PD protects against various DoS attacks, such as SYN flooding, SlowHTTP, and Memory DoS. The architecture dynamically adjusts security parameters, implements rate limiting, and executes appropriate mitigation strategies based on attack characteristics, ensuring continuous service availability.
III-F2 Enterprise Cloud Security Operations
In enterprise networks, LLM-PD enhances security operations by monitoring cloud infrastructure, automating alert triage, and implementing initial containment actions before human involvement. This significantly reduces detection and response times.
III-F3 Multi-Cloud Defense Orchestration
For organizations employing multi-cloud strategies, LLM-PD serves as a unified defense orchestrator that normalizes security telemetry across diverse cloud platforms. This enables coordinated defense actions across provider boundaries, preventing attackers from exploiting differences between cloud environments.
IV Case Study
IV-A Experimental Setups
It’s important to note that our LLM-PD architecture is model-agnostic and can be implemented with various LLMs. For our current implementation, we simulate a prototype of LLM-PD with GPT-4o mini due to its effective balance of reasoning capabilities and computational efficiency. Each agent is built through a series of prompt engineering techniques, and Table I presents the input prompts for each component.
To evaluate our proposed architecture in defeating threats, we implement a DoS attack scenario discussed in Section III-F. Specifically, for SYN flooding and SlowHTTP attacks, we simulate an elastic service that can create up to 10 replicas, using a total pool of 100 pods. Each pod supports 256 connections and a maximum memory utilization of 100. The initial setup activates 5 replicas, each comprising 10 pods. Additionally, we discuss Memory DoS, a co-resident attack that targets cloud infrastructure. Here we model a cluster with 5 racks, each containing 10 physical machines, and each machine capable of hosting up to 10 Virtual Machines (VMs). The maximum number of memory contention per VM is limited to 100 with runtime capped at 100 time steps. We conduct 10 episodes per test, in which an episode concludes when the system status remains stable for 5 consecutive steps.
We compare our proposed method with several well-known existing schemes, including DQN [6], Actor-Critic (AC) [13], and Proximal Policy Optimization (PPO) [14]. We adopt the same configurations and environments for training each scheme with the discount factor set to 0.98. The learning rate is 0.001 for the Q-network in DQN and the policy network in AC and PPO, and 0.01 for the value network in AC and PPO. All networks are shallow, with a single hidden layer containing 256 neurons. Since our method does not rely on training but rather on its own learning and evolution capabilities, these attacks are unseen by LLM-PD. To ensure statistical reliability and mitigate environmental uncertainties, we conduct 200 independent tests with initialization prior to each run.
IV-B Results
TABLE I: Prompts for each component in the proposed architecture.
Components Examples of Prompts
Basic Profile You are a security robot that can continuously improve security of cloud networks during multiple steps of attacks.
Collector You can call optional tools [Wireshark, Tcpdump, …] to collect data. Remove redundant information and output it to the [JSON, XML, …] format.
Analyzer You need to analyze the target cloud based on the current [Status] and evaluate the risk from [Indicators].
Decision-Maker Please decompose the task into subtasks based on [Priorities] and generate the optimal policy following [Constraints] and [Preferences]. The available defense actions are [MTD mechanisms, cyber deception methods, …].
Deployer The [Action] to be taken in this round is configured with [Configurations]. If it is not existed, please generate scripts and compile them for execution.
Feedback Giver Please evaluate the defense effectiveness based on [Security, Time, …] and the [Validity] of strategy with its implementation. Then, you need to store the entire process in [Memory] and feed it to other agents.
Figure 2: The surviving rate over 10 episodes when defending against different attacks with different strengths.
Fig. 2 presents the performance when defending against three attacks within different attack strengths, where the $x$-axis represents the number of episodes, the $y$-axis represents the average surviving rate under the current episode, and $n$ denotes the number of attackers. Note that the average survival rate is measured by the probability of successfully mitigating an attack at each time step after a defensive action is taken to protect the system. Additionally, an episode refers to a complete operation cycle of LLM-PD, and each episode retains the memory of feedback obtained from previous episodes. The results show that the average surviving rate gradually increases for all attacks and strengths when running tests for more episodes, which implies the effectiveness of the generated policy, the correct deployment of defense mechanisms, and especially, the significant increase brought by the learned experience. Although more attackers pose a challenge and slightly decrease its performance, our solution still exhibits high surviving rates of 88.8%, 92.1%, and 93.5% when defending against attacks launched from 50 attackers due to the guidance from the LLM-based agents.
We also present the number of average steps needed for successful defense against these attacks within 50 attackers in Table II, where $e$ denotes the number of episodes. There is an observation that the steps for mitigating SYN Flooding, SlowHTTP, and Memory DoS attacks are 9.79, 9.55, and 16.3, respectively, when $e=2$. This implies that the difficulty of defeating different attacks varies significantly when equipped with less knowledge and experience. As the episode progresses, the accumulation of knowledge allows our method to respond with fewer steps and reduced response times to successfully defend against various attacks. Specifically, the number of steps required converges to 7.51, 7.23, and 7.53 when $e=10$. This can be attributed to the fact that LLM-PD avoids ineffective strategies and actions based on previous memory, thus promoting efficient threat mitigation.
In addition, we compare LLM-PD with three previously discussed solutions in terms of success rate, and all methods are trained or prompted with the SYN Flooding scenario of 10 attackers in this comparison. It is worth noting that the success rate is a relatively broad indicator, focusing solely on whether the attack is ultimately resisted successfully, without considering whether all intermediate processes are optimal. As illustrated in Fig. 3, all methods exhibit more than 97.4% success rate, and especially, our solution achieves the highest performance when there are 10 attackers. Nevertheless, unlike existing approaches, LLM-PD encapsulates a vast spectrum of knowledge and a rich understanding from previous experience that enables it to act without additional or specific training. Not only that, when there are more attackers involved, LLM-PD exhibits remarkable adaptability, maintaining a high success rate while other approaches significantly reduce to 48.5%, 24.5%, and 31.5% when dealing with 50 attackers. This observation implies that LLM-PD can easily and swiftly adapt to new tasks or scenarios with relatively minimal additional training, offering a significant advantage over other methods that may require extensive retraining.
V Challenges and Future Directions
V-A Developing Context Management for Long-term Memory
The current implementation of LLM-PD faces limitations in maintaining contextual awareness across extended security incidents. While the architecture includes a sliding window mechanism, LLMs inherently struggle with managing long-term dependencies and correlating events across extended timeframes, which is crucial for detecting sophisticated advanced persistent threats (APTs) that unfold over weeks or months. Future work should explore specialized context management architectures for security applications, enabling more effective correlation of events across extended time periods. Developing temporal reasoning capabilities specific to attack progression patterns would significantly improve detection and mitigation of multi-stage, long-running security threats.
TABLE II: The number of average steps required to achieve successful defense under various attacks.
Attack Scenario Average Steps for Successful Defense
e = 2 e = 4 e = 6 e = 8 e = 10
SYN Flooding 9.79 8.47 7.60 7.55 7.51
SlowHTTP 9.55 8.62 7.72 7.75 7.23
Memory DoS 16.3 11.6 8.50 7.89 7.53
Figure 3: The comparison of success rate against different numbers of attackers when equipped with different defensive schemes.
V-B Designing Resource-Efficient LLM Based Security Solution
LLM-PD faces practical deployment challenges related to computational resource demands and integration with existing security infrastructure. High-capability LLMs require significant computational resources, which can introduce latency in time-sensitive scenarios and limit deployment in resource-constrained environments. Future efforts should focus on optimizing LLM-PD for resource efficiency through techniques like model distillation and specialized model architectures that maintain reasoning capabilities while reducing resource requirements. These advancements would transform LLM-PD from a promising research direction into a practical, widely deployable security solution for diverse cloud environments.
V-C Improving Robustness against Potential Adversarial Inputs
Another significant challenge specific to LLM-PD is its potential vulnerability to adversarial inputs designed to manipulate the defense system. Attackers familiar with LLMs might craft network traffic patterns or system behaviors specifically designed to trigger false interpretations or ineffective defense responses. To counter this, future research should focus on enabling LLM-PD to recognize and resist such manipulation attempts. This includes implementing robust verification mechanisms that cross-check LLM interpretations against multiple independent detection systems and developing techniques to identify inconsistencies between observed attack patterns and known attack taxonomies.
VI Conclusion
The development of LLMs is promising for tackling challenges associated with mitigating cyberattacks. In this paper, we first introduce an innovative proactive defense architecture empowered by LLM to facilitate security improvement and threat mitigation within the cloud environment. By leveraging the advanced capabilities of LLM in data collection, security analysis, task inference, defense deployment, and effectiveness evaluation, our method is capable of thoroughly analyzing the security situation, efficiently executing suitable actions, and continuously evolving itself to adapt to varying and complex attack scenarios. Then, a detailed case study on three kinds of DoS attacks is discussed. Experimental results demonstrate that the proposed architecture improves effectiveness and efficiency of defense compared to the state-of-the-art methods. Finally, we highlight future directions and open challenges to further enhance cloud security.
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grant No. 62202097, in part by China Postdoctoral Science Foundation under Grant No. 2024T170143 and No. 2022M710677, and in part by Jiangsu Funding Program for Excellent Postdoctoral Talent under Grant No. 2022ZB137.
Biographies
Yuyang Zhou [M’22] (yyzhou@seu.edu.cn) is currently working as a postdoc with the School of Cyber Science and Engineering, Southeast University.
Guang Cheng [M’04] (chengguang@seu.edu.cn) is a Full Professor in the School of Cyber Science and Engineering, Southeast University.
Kang Du (220235194@seu.edu.cn) is currently pursuing his master degree with the School of Cyber Science and Engineering at Southeast University.
Zihan Chen [M’23] (zhchen@njnet.edu.cn) is a postdoc researcher with the School of Cyber Science and Engineering at Southeast University.
Yuyu Zhao [M’23] (yyzhao@seu.edu.cn) is a lecturer with the School of Cyber Science and Engineering at Southeast University.